OpenStudy (anonymous):

What are the x- and y-intercepts of the line tangent to the circle (x - 2)^2 + (y - 2)^2 = 25 at the point (5,6)?

OpenStudy (anonymous):

Implicit differentiate to get \[ 2 ( x-2) + 2 (y-2) y'=0\\ y'= \frac{2-x}{y-2}\\ \] Find y' when x=5 and y=6

OpenStudy (anonymous):

\[ y'=m = \frac{ 2 -5}{6-2}= -\frac 3 4 \]

OpenStudy (anonymous):

\[ y- 6 = -\frac 3 4 ( x-5) \]

OpenStudy (anonymous):

See the circle and the line on the atteched graph.

OpenStudy (anonymous):

\[ y- 6 = -\frac 3 4 ( x-5)\\ y = -\frac 3 4 ( x-5) + 6 \] If x=0 then \[ y=\frac {39}4 \] if y=0 \[ x=13 \]